# Confidence-Based Annotation Of Brain Tumours In Ultrasound

### v1

@@ 4 Discussion
A clinical imaging perspective of Eq. 2 is provided referencing standard practices in radiology. Fully informed diagnostic decision making with MRI requires multiple sequencing. As stated in [27], reliance on individual sequences such as T1-weighted or T2-weighted, will underestimate the spread of cancer in the surrounding areas of the determined region of interest. For example, [28] measured that the use of proton magnetic resonance spectroscopic imaging, on average, identified tumour margin extensions greater than $20\%$ from what was identifiable using T2 weighted images. We extrapolate this with the case of standard B-mode US. An example of this can be seen in the use of superb microvascular imaging (SMI). In [29], it was seen that the use of SMI, particularly for high-grade gliomas, improved the visualisation of the tumour margin and was associated with an increased likelihood of gross total resection. This clinical perspective further emphasises the limitation of B-mode segmentation and the issue of discretely determining the edge point. The evidence of margin difference when using multiple modalities or sequences indicates not only that discrete margin delineation is not achievable purely using B-mode imaging. However, even with multiple sources of information, the margin uncertainty is only reduced, not eliminated, because of the infiltrative properties.
@@ 5 Conclusion
The challenge of discretely annotating diffuse, infiltrative, brain tumours in B-mode US is demonstrated. A logic framework is presented to formally construct the problem specific to tumour margin aleatoric uncertainty. A method for confidence-based annotation is proposed with a protocol derived from computer vision and radiology theory. The method enables sparse and fast annotation which should facilitate rapid dataset generation. The confidence based protocol reduces and/or removes the need for averaging interobserver annotations - through minimising the annotator level epistemic uncertainty by the design of the annotation protocol and method. Evaluation is performed to measure how well the protocol and KDE process can, implicitly, estimate the interobserver variance of 4 professionals, discretely annotating the corresponding images. A reliability diagram is created and an average bin accuracy of 0.1 was measured which indicates correlation between the confidence method and the interobserver variance. In particular, a linear relationship was measured along the tumour margin. With a Pearson correlation of 0.8 and Spearman correlation of 0.75. The downstream application of the confidence annotations is investigated. A UNet is trained for segmentation on both the neuroradiologist discrete annotations (hard label) and the confidence annotations (soft label). Across all evaluation folds, the soft labels produced a better calibrated network with better Brier scores.
The limitations of this study include the potential bias in the analysis of the relationship between confidence annotation and interobserver variation. Although neurosurgeon annotations were not seen during the confidence annotation process, the same authors participated in the prior study, which may introduce bias due to the influence of memory. However, this bias was controlled through strict adherence to the annotation protocol and by anchoring the localisation process on top of the neuroradiologist’s annotations.
Future work to remove interposed normal tissue from the confidence map could be explored, perhaps using sink points.

### v2

@@ 4 Discussion
A clinical imaging perspective of Eq. LABEL:eq:fuzzybound is provided referencing standard practices in radiology. Fully informed diagnostic decision making with MRI requires multiple sequencing. As stated in [27], reliance on individual sequences such as T1-weighted or T2-weighted, will underestimate the spread of cancer in the surrounding areas of the determined region of interest. For example, the work in [28] measured that the use of proton magnetic resonance spectroscopic imaging, on average, identified tumour margin extensions greater than $20\%$ from what was identifiable using T2 weighted images. We extrapolate this with the case of standard B-mode US. An example of this can be seen in the use of superb microvascular imaging (SMI). In [29], it was seen that the use of SMI, particularly for high-grade gliomas, improved the visualisation of the tumour margin and was associated with an increased likelihood of gross total resection. This clinical perspective further emphasises the limitation of B-mode segmentation and the issue of discretely determining the edge point. The evidence of margin difference when using multiple modalities or sequences indicates not only that discrete margin delineation is not achievable purely using B-mode imaging. However, even with multiple sources of information, the margin uncertainty is only reduced, not eliminated, because of the infiltrative properties.
@@ 5 Conclusion
The challenge of discretely annotating diffuse, infiltrative, brain tumours in B-mode US is demonstrated. A method for confidence-based annotation is proposed with a protocol derived from computer vision and radiology theory. The method enables sparse and fast annotation which should facilitate rapid dataset generation. The confidence based protocol reduces and/or removes the need for averaging interobserver annotations - through minimising the annotator level epistemic uncertainty by the design of the annotation protocol and method. Evaluation is performed to measure how well the protocol and KDE process can, implicitly, estimate the interobserver variance of 4 professionals, discretely annotating the corresponding images. A reliability diagram is created and an average bin accuracy of 0.1 was measured which indicates correlation between the confidence method and the interobserver variance. In particular, a linear relationship was measured along the tumour margin, with a Pearson correlation of 0.8 and Spearman correlation of 0.75. The downstream application of the confidence annotations is investigated. A UNet is trained for segmentation on both the neuroradiologist discrete annotations (hard label) and the confidence annotations (soft label). Across all evaluation folds, the soft labels produced a better calibrated network with better Brier scores.
The limitations of this study include the potential bias in the analysis of the relationship between confidence annotation and interobserver variation. Although neurosurgeon annotations were not seen during the confidence annotation process, the same authors participated in the prior study [12], which may introduce bias due to the influence of memory. However, this bias was controlled through strict adherence to the annotation protocol and by anchoring the localisation process on top of the neuroradiologist’s annotations. Future work to remove interposed normal tissue from the confidence map could be explored, perhaps using sink points.
@@
Code, Data, and Materials Availability
The authors do not have permission to share the data set used in this study. The code can be made available upon reasonable request to the corresponding author.